OpenStudy (anonymous):

The function f(x)=245(4)^x represents the growth of a fruit fly population every YEAR in the orange grove. Stacy wants to manipulate the formula to an equivalent form that calculates every MONTH, not every year. Which function is correct for Stacy's purposes? f(x)=45(4)^x f(x)=45(4^12)x/12 f(x)=2752(4)^x f(x)=245(4^1/12)^12x II think that it is the second option but idk. My reasoning is because since in a single year there are 12 months then it should be the denominator of a fraction and the numerator would be howm many months. But im also not sure if it might be the last one

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

@campbell_st

OpenStudy (mathmath333):

i think ur options are not in proper format, try to write it exactly again

OpenStudy (anonymous):

The function f(x)=245(4)^x represents the growth of a fruit fly population every YEAR in the orange grove. Stacy wants to manipulate the formula to an equivalent form that calculates every MONTH, not every year. Which function is correct for Stacy's purposes? f(x)=45(4)^x f(x)=45(4^12)x/12 f(x)=2752(4)^x f(x)=245(4^1/12)^12x II think that it is the second option but idk. My reasoning is because since in a single year there are 12 months then it should be the denominator of a fraction and the numerator would be howm many months. But im also not sure if it might be the last one

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

\(\large f(x)=45(4)^x\\ \large f(x)=45(4^{12})*\frac{x}{12}\\ \large f(x)=2752(4)^x \\ \large f(x)=245(4^{\frac{1}{12}})^{12x }\) are this correct ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

@CoconutJJ any help?

OpenStudy (anonymous):

@Abhisar

OpenStudy (anonymous):

#@surryyy

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{f(x)=245(4)^x \hspace{.33em}\\~\\ f(1)=245(4)^{1}=980\hspace{.33em}\\~\\ \normalsize \text{option A} \hspace{.33em}\\~\\ 45(4)^1=180 \hspace{.33em}\\~\\ \normalsize \text{option B} \hspace{.33em}\\~\\ 45(4^{12})*\frac{1}{12}\approx 12 million \hspace{.33em}\\~\\ \normalsize \text{option C} \hspace{.33em}\\~\\ 2752(4)^x =11008 \hspace{.33em}\\~\\ \normalsize \text{option D} \hspace{.33em}\\~\\ 245(4^{\frac{1}{12}})^{12}=980 \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

@mathmath333 the question isnt asking to solve theses, its asking which one would she use if she wanted to use it in months instead of years

OpenStudy (mathmath333):

i know i showed all options

OpenStudy (mathmath333):

of which D mathches with ur function , but m not sure.

OpenStudy (anonymous):

@mathmath333 thanks alot , i think ive got it

OpenStudy (mathmath333):

D is correct ?

OpenStudy (anonymous):

@mathmath333 i believe so lol